OpenStudy (anonymous):

help! A polynomial function can be written as (x - 1)(x - 4)(x + 7). What are the x-intercepts of the graph of this function? a. (1, 0), (4, 0), (7, 0) b. (-1, 0), (-4, 0), (-7, 0) c. (1, 0), (4, 0), (-7, 0) d. (-1, 0), (-4, 0), (7, 0)

OpenStudy (anonymous):

Set it equal to 0. What do you know about quadratic equations in their factored form that allow you to get the roots? How is that applied here?

OpenStudy (ikram002p):

to find the x-intercepts that mean y = 0 y=(x - 1)(x - 4)(x + 7) 0=(x - 1)(x - 4)(x + 7) so we conclude that on of the multible should be 0 0= x - 1 or 0= x - 4 or 0= x +7 find x values in the three cases.